Please help
Solve each equation. Round your answers to the nearest hundreth, if necessary.
1. 8(x-7)^2 = 200
2. 4/3(x^2 +12)- 25 = 11
Thank you very much for your willingness to help me.

Answer by fractalier(6550)

From 8(x-7)^2 = 200, we divide both sides by 8 and get
(x-7)^2 = 25 Now take the square root of both sides
x - 7 = ±5 so that
x = 7 ± 5 = 12 or 2
From 4/3(x^2 + 12) - 25 = 11, we add 25 to both sides, then multiply by 3/4 and get
4/3(x^2 + 12) = 36
x^2 + 12 = 27
From here, subtract 12 and take the root and get
x = ± radical 15